# Discontinuous Transition to Active Nematic Turbulence

### v1

@@ IV.3 Second subcritical bifurcation and the onset of chaotic transients
The oscillations that emerged in the previous bifurcation persist through the transition to elongated vortices (Fig. 2). Thus, we suggest that this transition corresponds to a subcritical Hopf bifurcation 61. Consistent with this scenario, we observe growing oscillations after the three-vortex state becomes unstable (early stage of Fig. 2). These oscillations then become transiently chaotic as the system explores a vast number of possible configurations in the search for a new stable state (middle stage of Fig. 2). The flow rapidly switches between different patterns, including brief visits to almost stable patterns, until it finally spirals down into an elongated vortex state (final stage of Fig. 2, Movie S3). Overall, the second subcritical bifurcation yields the earliest (transient) appearance of chaos in active nematics. By bringing about bistability, oscillations, and chaos, the two subcritical bifurcations that we found provide the key dynamical ingredients for the transition to chaos at higher activity discussed in Section III.
@@
Chaotic transients appear with the elongated vortices state, which exists above $A\approx 1373$. This first appearance of chaos can also be understood from the statistical approach of Section III. Fig. 1 shows that, in the chaotic state, the MLE is positive and increases with activity. We ask: What is the smallest activity that allows for a positive MLE? To answer this question, we fit a line to the positive MLE values in Fig. 1 and we extrapolate it to smaller activities (Fig. S2). This extrapolated line could indicate the MLE of the short chaotic transients at $A\lesssim 3800$. The extrapolation crosses zero at $A=1380$ (Fig. S2), which is consistent with our finding that chaotic transients first appear at $A\approx 1373$.
IV.4 Towards higher activities
@@
As activity increases even further, more and more flow states appear and disappear through a variety of bifurcations. The state space becomes increasingly complex, with many coexisting stable attractors. A hint of this complexity is already given by the fact that, as activity decreases, the elongated-vortex state suffers an intermediate destabilization (kink around $A=1430$ in Fig. 2) before transitioning to the three-vortex state. The plethora of possible flow states at these activities includes complex oscillations such as those at $A=1560$ shown in Movie S1. Sampling the state space beyond the range of activities covered in Fig. 2 is an interesting direction for future work.
V Discussion and conclusions
@@
In summary, we have found a discontinuous transition to unconfined active nematic turbulence. This result contrasts with the continuous transition found both for active nematics in channel confinement 47 as well as for high-Reynolds pipe flow 3; 4; 5; 6; 7; 8; 9; 10; 11; 12. These transitions have critical behavior in the directed percolation universality class. Our finding of a discontinuous transition reveals that the onset of active nematic turbulence needs not exhibit universal behavior.
@@
Why does the transition change from continuous to discontinuous in the absence of confinement? We propose that the transition is discontinuous due to the long-ranged hydrodynamic interactions of Stokes flow. Through them, the flow field is coupled across the entire system, and hence it is either globally laminar or globally chaotic. Indeed, our simulations never show spatial coexistence of laminar and turbulent regions. As a result, the system has to switch discontinuously to chaos.
@@
In contrast, the directed-percolation transition takes place through spatiotemporal intermittency, in which localized puffs of turbulence coexist with a laminar background 11. This spatial coexistence allows the turbulent phase to emerge continuously: As either the activity or the Reynolds number increases, the turbulent phase fills more space. We speculate that spatial coexistence of different flow states is possible for confined active nematics because the no-slip condition at the channel walls screens the long-ranged hydrodynamic interactions, thus decoupling different regions of the system 47; 52. Similarly, in high-Reynolds flows, flow perturbations do not propagate instantaneously across the system, which enables turbulent puffs to progressively invade laminar regions in the process of spatiotemporal intermittency 3; 11. Recent work showed that suppressing spatiotemporal intermittency, either by numerically eliminating turbulent structures 64 or via body forces in experiments 13, renders the transition to inertial turbulence discontinuous. Analogously, we propose that, in unconfined active nematics, hydrodynamic interactions suppress spatiotemporal intermittency, hence yielding a discontinuous transition.
@@
We propose to test our prediction of a discontinuous transition to turbulence in experiments of active nematics without boundaries, which can be realized by placing microtubule-kinesin mixtures on toroidal droplets 65. In such experiments, activity could be controlled using light-activated motors 66; 67, and the discontinuous nature of the transition could be revealed, for example, by the presence of hysteresis as shown in Fig. 1.
@@
Beyond the nature of the transition to turbulence, our study also probed the dynamical route towards it. We found that, after the well-known spontaneous-flow instability, active nematics experience subcritical bifurcations whereby they discontinuously switch to different laminar flow states (Fig. 2). Upon these bifurcations, chaotic transients emerge at activities $A\gtrsim 1373$, much below the transition to turbulence at $A\approx 4900$. In dynamical-systems terms, such chaotic transients are characterized as a chaotic saddle — a region of phase space where the system explores many unstable configurations before exiting the saddle when it finds a stable state 68; 10. At low activities, this search is quickly successful, and the chaotic transients are short. As activity increases, stable states become harder to find, resulting in longer transients (Fig. S1). When stable states are practically not possible to find, chaotic transients become arbitrarily long, resulting in sustained active turbulence. Our results suggest that this picture, which was previously put forward for inertial turbulence 68; 10, also applies to active nematic turbulence.
@@
We obtained our results using a minimal model of active nematics, which features neither flow alignment nor topological defects 55. Future work could address the role that these ingredients have on the transition to active turbulence. Moreover, controlling the transition to turbulence is relevant for biological systems such as bacterial colonies or epithelial monolayers 69, which are often in contact with a substrate. Therefore, future work should also investigate how substrate friction affects the results presented here.
@@
To conclude, our work provides new pieces of the connection between active and inertial turbulence. Whereas most work so far focused on the statistics of the chaotic flow 15, here we revealed that the route to active turbulence also displays key similarities and differences with its high-Reynolds counterpart.
@@ Acknowledgments
We thank Marc Avila, Jaume Casademunt, Björn Hof, Jean-François Joanny, and Holger Kantz for discussions. We acknowledge computing support by the Max Planck Computing and Data Facility and the computing facility at MPI-PKS. RA thanks the KITP Program Active Solids: From Metamaterials to Biological Tissue, where work on this paper was undertaken. This research was supported in part by grant no. NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).
Appendix A Methods
A.1 Numerical Implementation
Following Ref.55, we numerically solve the force balance Eq. 1 spectrally using $256\times 256$ Fourier modes and the $2/3$ anti-aliasing rule. We evolve the director dynamics Eq. 4 using an alternating direction implicit (ADI) method with finite differences 70, on a grid of $256\times 256$ points with a time step of $\Delta t=1.8\times 10^{-5}\tau_{\mathrm{r}}$, parallelized for GPU computation using CUDA. Unless otherwise stated for initial transients, the simulations are deterministic, without additional random noise.
Also following Ref. 55, the simulations for Figs. 1, 1 and 1 include additive Gaussian white noise with amplitude $D=5\times 10^{-4}L^{2}/\tau_{\mathrm{r}}$ on the director field for an initial transient of $10^{-2}\tau_{\mathrm{r}}$, after which integration continues without noise up to a maximum time of $6\times 10^{2}\tau_{\mathrm{r}}$. A shorter integration time of $10\tau_{\mathrm{r}}$ was used for activities below $A=3000$, as in every case the flow reached a (statistically) steady state already long before this time. Figs. 1, 1 and 1 show results using 32 independent realizations for cases with $A\leq 3700$, and 64 realizations from the onset of the transition region $A\geq 3800$. For the simulations in Fig. 1, we start from an aligned nematic with $\theta(\bm{r})=0$ at $A=0$, and we then increase activity in steps of 100, allowing $2.5\tau_{r}$ time units for the system to relax to the new state before increasing activity further. Hysteresis is demonstrated by taking the final state as an initial condition and decreasing activity in the same way.
The results of Fig. 2 come from simulations starting with an aligned nematic with $\theta(\bm{r})=0$ at $A=1$. Then, every $2\tau_{\mathrm{r}}$ time units the activity is increased by one, e.g. $A=1\rightarrow 2$. Below $A=100$, an initial burst of white noise with very small amplitude ($D=5\times 10^{-6}L^{2}/\tau_{\text{r}}$) is added each time the activity is increased to speed up the search of the new state. The final values of $\langle v^{2}\rangle_{\mathbf{r}}$, $\langle v_{x}^{2}\rangle_{\mathbf{r}}$, and $\langle v_{y}^{2}\rangle_{\mathbf{r}}$, as shown in Fig. 2, are recorded before the activity is increased again. The regions showing multistability were probed by taking the state above the bifurcation point as an initial condition, and then following the same protocol but with decreasing activity. Lyapunov exponent calculations were carried out using the final state as an initial condition, and a final integration time of $2\times 10^{2}\tau_{\mathrm{r}}$ was used to ensure convergence of the deviation vectors.
A.2 Lyapunov Exponents
In principle, for a dynamical system given by $\dot{\bm{x}}=\bm{f}(\bm{x})$, the maximal Lyapunov exponent (MLE) $\Lambda$ of a state $\bm{x}_{0}$ is defined as 71; 72; 73
$\Lambda(\bm{x}_{0})=\lim_{t\rightarrow\infty}\frac{1}{t}\log\left(\frac{||\bm{w}(t)||}{||\bm{w}(0)||}\right),$ (6)
where $\bm{w}(t)$ is an infinitesimal perturbation to the state $\bm{x}_{0}$. Equation 6 shows that a subexponential growth of the deviation vector $\bm{w}(t)$ will yield a value of zero for the MLE, indicating non-chaotic dynamics without sensitive dependence on initial conditions. In contrast, exponential growth in $\bm{w}(t)$ will result in a non-zero value for $\Lambda$. The faster the deviation grows, the larger $\Lambda$ will be, thus capturing the intuitive notion of it being more sensitively dependence on initial conditions.
In practice, the infinite time limit in Eq. 6 is replaced by a finite time that is long enough to capture the dynamics of the system. Similarly, in chaotic cases, the exponential growth of $\bm{w}(t)$ requires us to periodically rescale it (see Ref. 73 and references therein). Here, we rescaled the deviation back to a norm of $||\bm{w}||=10^{-6}$ every $1\times 10^{-3}\tau_{\mathrm{r}}$. Varying the parameters of this rescaling had no bearing on the results.
For the active nematic model considered here, after discretizing the time and space derivatives, the dynamical system is given by $\dot{\theta}_{ij}=f_{ij}(\theta)$, where $i$ and $j$ indicate the discrete coordinates of the grid points. Thus, we define the deviation as $w_{ij}=\tilde{\theta}_{ij}-\theta_{ij}$, where $\tilde{\theta}$ is a field given by adding to $\theta$ a small random perturbation drawn from a normal distribution. We then compute $\bm{w}(t)$ by integrating both $\theta$ and $\tilde{\theta}$ in time and using the finite-time version of Eq. 6.
The spectrum of LEs is computed through Gram-Schmidt orthonormalization of several different deviation vectors $\bm{w}^{(m)}$ 73 (one deviation vector per LE), with these deviations to the reference $\theta$ field obtained in the same way as for the MLE. Due to the absence of spatially coexisting regular and chaotic dynamics, considering the growth of the global norm of the deviation provides complete information about the overall chaotic dynamics of the system.
Finite time LEs or stretching numbers $\alpha$ are computed as the growth rate of the deviation vector over a finite time,
$\alpha_{k}=\frac{1}{t_{k}-t_{k-1}}\log\left(\frac{||\bm{w}(t_{k})||}{||\bm{w}(t_{k-1})||}\right),$ (7)
where $k$ labels the time window from $t_{k-1}$ until $t_{k}$. Here, we rescaled the deviation vector every $1\times 10^{-3}\tau_{\mathrm{r}}$, so that $||\bm{w}(t_{k-1})||=||\bm{w}(0)||$. We record stretching numbers before every rescaling. The stretching numbers $\alpha$ used in the results of Figs. 1 to 1 are computed over time windows of $1\times 10^{-2}\tau_{\mathrm{r}}$.
@@ A.2 Lyapunov Exponents
[…]
Figure S1: Lifetimes of chaotic transients. a, The probability $P_{\mathrm{t}}$ of a chaotic transient surviving a certain time $t$ decays exponentially for different values of the activity number $A$. The dashed lines indicate a fit to an exponential function $P_{\mathrm{t}}=\exp(-t/\tau_{\mathrm{t}})$, from which we obtain the lifetime $\tau_{\mathrm{t}}$. b, The mean lifetime $\tau_{\mathrm{t}}$ of chaotic transients increases very sharply (by orders of magnitude) as the activity number $A$ approaches the region of the transition to turbulence, which begins around $A\approx 3800$. The unit of time in this figure is $\tau_{\mathrm{r}}$, as defined in the Main Text.

### v2

@@ Abstract
The oscillations that emerged in the previous bifurcation persist through the transition to elongated vortices (Fig. 2). Thus, we suggest that this transition corresponds to a subcritical Hopf bifurcation 68. Consistent with this scenario, we observe growing oscillations after the three-vortex state becomes unstable (early stage of Fig. 2). These oscillations then become transiently chaotic as the system explores a vast number of possible configurations in the search for a new stable state (middle stage of Fig. 2). The flow rapidly switches between different patterns, including brief visits to almost stable patterns, until it finally spirals down into an elongated vortex state (final stage of Fig. 2, Movie S3). Overall, the second subcritical bifurcation yields the earliest (transient) appearance of chaos in active nematics. By bringing about bistability, oscillations, and chaos, the two subcritical bifurcations that we found provide the key dynamical ingredients for the transition to chaos at higher activity shown in Fig. 1.
@@
Chaotic transients appear with the elongated vortices state, which exists above $A\approx 1373$. This first appearance of chaos can also be understood from the statistical approach of the previous section. Fig. 1 shows that, in the chaotic state, the MLE is positive and increases with activity. We ask: What is the smallest activity that allows for a positive MLE? To answer this question, we fit a line to the positive MLE values in Fig. 1 and we extrapolate it to smaller activities (Fig. S3). This extrapolated line could indicate the MLE of the short chaotic transients at $A\lesssim 3800$. The extrapolation crosses zero at $A=1380$ (Fig. S3), which is consistent with our finding that chaotic transients first appear at $A\approx 1373$.
Towards higher activities
@@
As activity increases even further, more and more flow states appear and disappear through a variety of bifurcations. The state space becomes increasingly complex, with many coexisting stable attractors. As an illustration, Fig. S4 shows four possible vortex states found at $A=2400$. A hint of this complexity is already given by the fact that, as activity decreases, the elongated-vortex state suffers an intermediate destabilization (kink around $A=1430$ in Fig. 2) before transitioning to the three-vortex state. The plethora of possible flow states at these activities includes complex oscillations such as those at $A=1560$ shown in Movie S1. Sampling the state space beyond the range of activities covered in Fig. 2 is an interesting direction for future work.
Discussion
Differences with directed percolation
@@ Discussion
In summary, we have found a discontinuous transition to unconfined active nematic turbulence. This result contrasts with the continuous transition found both for active nematics in channel confinement 51 as well as for high-Reynolds pipe flow 3, 4, 5, 6, 7, 8, 9, 10, 11, 12. These transitions have critical behavior in the directed percolation universality class. Our finding of a discontinuous transition reveals that the onset of active nematic turbulence needs not exhibit universal behavior.
@@
Why does the transition change from continuous to discontinuous in the absence of confinement? We propose that the transition is discontinuous due to the long-ranged hydrodynamic interactions of Stokes flow. Through them, the flow field is coupled across the entire system, and hence it is either globally laminar or globally chaotic. To probe this idea, we followed Ref. 51 and obtained kymographs of enstrophy, $\epsilon(x,t)=|\omega(x,t)|^{2}$, where $\omega$ is the vorticity. Upon an increase in activity, we find a global switch from laminar oscillatory flow to chaos (Fig. 3). All points in the system become chaotic simultaneously; we find no spatial coexistence of laminar and turbulent regions. As a result, the system has to switch discontinuously to chaos.
[…]
Figure 3: Global transition to turbulence without spatial coexistence of laminar and chaotic domains. Kymograph of the enstrophy $\epsilon(x,t)=\langle\epsilon(x,y,t)\rangle_{y}$ averaged over the $y$ axis (3) and time trace at position $x=0.33$ (3). The activity is increased from $A=4850$ to $A=4900$ at $t=0.3\tau_{\text{r}}$. As a result, the flow transitions from a stable oscillating vortex to unstable oscillations and aperiodic motion. The transition to chaos at $t\approx 0.87\tau_{\text{r}}$ is temporally sudden and spatially global, unlike the localized turbulent puffs characteristic of the directed percolation scenario.
@@
In contrast, the directed-percolation transition takes place through spatiotemporal intermittency, in which localized puffs of turbulence coexist with a laminar background 11. This spatial coexistence allows the turbulent phase to emerge continuously: As either the activity or the Reynolds number increases, the turbulent phase fills more space. We speculate that spatial coexistence of different flow states is possible for confined active nematics because the no-slip condition at the channel walls cuts off the long-ranged hydrodynamic interactions, thus decoupling different regions of the system 51, 56. Similarly, in high-Reynolds flows, flow perturbations do not propagate instantaneously across the system, which enables turbulent puffs to progressively invade laminar regions in the process of spatiotemporal intermittency 3, 11. Recent work showed that suppressing spatiotemporal intermittency, either by numerically eliminating turbulent structures 70 or via body forces in experiments 13 and theory 14, renders the transition to inertial turbulence discontinuous. Analogously, we propose that, in unconfined active nematics, hydrodynamic interactions suppress spatiotemporal intermittency, hence yielding a discontinuous transition.
Role of system size
In the directed-percolation scenario, the system must be large enough to allow for the spatial coexistence of laminar and turbulent patches. Here, however, the system size $L$ enters the definition of the dimensionless activity number $A=L^{2}/\ell_{\text{c}}^{2}$, where $\ell_{\text{c}}$ is the active length. Therefore, the system size cannot be varied without also affecting the dimensionless activity. This fact is a physical feature of active nematics: Because of the long-ranged hydrodynamic interactions of Stokes flow, the spontaneous-flow instability is long-wavelength. Thus, increasing active stresses (decreasing the active length) is equivalent to increasing the system size 67, 71, 59; both variations result in more unstable modes and can thus trigger the transition to chaos. This feature makes conventional finite-size analyses impossible in our setting. To circumvent this issue, we instead performed a finite-time analysis (Fig. 1), which revealed that the transition to turbulence takes place at $A^{*}\approx 4900$. Below it, long chaotic transients end up laminarizing; above it, chaos persists.
Proposal of an experimental test
@@
We propose to test our prediction of a discontinuous transition to turbulence in experiments of active nematics without boundaries, which can be realized by placing microtubule-kinesin mixtures on toroidal droplets 72. In such experiments, activity could be controlled using light-activated motors 73, 74, and the discontinuous nature of the transition could be revealed, for example, by the presence of hysteresis as shown in Fig. 1.
Other discontinuous transitions in active turbulence
We now compare our results to discontinuous transitions found in active turbulence. In a model where activity enters as a scale-dependent effective viscosity, Linkmann et al. found a discontinuous transition between regimes of active and hydrodynamic turbulence 75. Here, in contrast, we report a discontinuous transition between laminar and chaotic flow. Respectively, in the Toner-Tu-Swift-Hohenberg model of active polar fluids, James et al. and Reinken et al. found a discontinuous transition from a vortex lattice to turbulence 43, 45. Unlike in our case, however, that transition can be triggered by two different activity-related parameters, and the two states can coexist in space. Similarly, recent experiments with microtubule-based active nematics showed the spatial coexistence of turbulent and laminar domains triggered by the nematic-to-smectic transition of a surrounding molecular liquid crystal 76. These findings highlight the potential of using surrounding fluids to control the transition to turbulence in experiments.
The chaotic saddle
@@
Beyond the nature of the transition to turbulence, our study also probed the dynamical route towards it. We found that, after the well-known spontaneous-flow instability, active nematics experience subcritical bifurcations whereby they discontinuously switch to different laminar flow states (Fig. 2). Upon these bifurcations, chaotic transients emerge at activities $A\gtrsim 1373$, much below the transition to turbulence at $A\approx 4900$. In dynamical-systems terms, such chaotic transients are characterized as a chaotic saddle — a region of phase space where the system explores many unstable configurations before exiting the saddle when it finds a stable state 77, 10. At low activities, this search is quickly successful, and the chaotic transients are short. As activity increases, stable states become harder to find, resulting in longer transients (Fig. S1). When stable states are practically not possible to find, chaotic transients become arbitrarily long, resulting in sustained active turbulence. Our results suggest that this picture, which was previously put forward for inertial turbulence 77, 10, also applies to active nematic turbulence.
The vortex-packing hypothesis
In the transition region, we found an intriguing V-shaped behavior of the chaotic fraction (Fig. 1). Whereas more work is required to understand it, here we speculate about a potential mechanism. As activity increases, the active length decreases; hence, vortices become smaller, which allows more of them to fit into the system. Since vortices are discrete objects, there could be specific activity values at which they pack particularly well into the system. These vortex patterns could be easily-accessible stable attractors, which could cause the decrease in the chaotic fraction in the middle of the transition region, before persistent chaos finally sets in. We defer an exploration of this vortex-packing hypothesis to future work.
Generality of the results
We obtained our results using a minimal model of active nematics, which features neither flow alignment nor topological defects 59. Our work suggests that the discontinuous nature of the transition is due to the long-ranged hydrodynamic interactions of Stokes flow, which exist also in the presence of flow alignment and topological defects. Thus, we expect that these ingredients will not change the nature of the transition. Future work could test this expectation in simulations including defects and flow alignment. In particular, adding flow alignment makes systems with extensile and contractile stresses no longer equivalent, and hence future work could test if the sign of active stresses impacts the transition to turbulence.
@@
Beyond the nature of the transition, both defects 57, 58, 54, 26, 78, 79 and flow alignment 61, 64 strongly affect the flow patterns, and hence they could impact the sequence of bifurcations leading to chaos (Fig. 2). Thus, our results provide a basis for future work to address the role of defects and flow alignment on the transition to active turbulence by exposing which additional features they bring about. Moreover, future work should also investigate how our results are affected by substrate friction, which is relevant for biological systems such as bacterial colonies or epithelial monolayers 80.
Conclusion
@@
To conclude, our work provides new pieces of the connection between active and inertial turbulence. Whereas most work so far focused on the statistics of the chaotic flow 16, here we revealed that the route to active turbulence also displays key similarities and differences with its high-Reynolds counterpart.
@@ Acknowledgments
We thank Marc Avila, Jaume Casademunt, Amin Doostmohammadi, Peter Hampshire, Björn Hof, Jean-François Joanny, Holger Kantz, and Kristian Thijssen for discussions. We thank Ekaterina Maximova for related preliminary work. We acknowledge computing support by the Max Planck Computing and Data Facility and the computing facility at MPI-PKS. RA thanks the KITP Program Active Solids: From Metamaterials to Biological Tissue, where work on this paper was undertaken. This research was supported in part by grant no. NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).
Author contributions
M.H. performed and analyzed the simulations. R.A. conceived and supervised the work. M.H. and R.A. discussed the results and wrote the paper.
Competing interests
The authors declare no competing interests.
Data availability statement
The data produced for this work is available at https://gitlab.pks.mpg.de/hillebrand/transition-to-active-turbulence.
Code availability statement
The codes to produce the simulations in this work are available at https://gitlab.pks.mpg.de/hillebrand/transition-to-active-turbulence.
@@
Figure S1: Lifetimes of chaotic transients. a, The probability $P_{\mathrm{t}}$ of a chaotic transient surviving a certain time $t$ decays exponentially for different values of the activity number AA. The dashed lines indicate a fit to an exponential function $P_{\mathrm{t}}=\exp(-t/\tau_{\mathrm{t}})$, from which we obtain the lifetime $\tau_{\mathrm{t}}$. b, The mean lifetime $\tau_{\mathrm{t}}$ of chaotic transients increases very sharply (by orders of magnitude) as the activity number AA approaches the region of the transition to turbulence, which begins around $A\approx 3800$. The unit of time in this figure is $\tau_{\mathrm{r}}$, as defined in the Main Text.
Figure S2: The spectrum of Lyapunov exponents reveals a sharp transition to high-dimensional chaos. The largest six Lyapunov exponents measured after a time $t_{\text{f}}=5000\tau_{\text{r}}$ for a single realization. Below the critical activity $A^{*}\approx 4900$, all Lyapunov exponents are zero or negative. As the stable laminar state approaches the transition to chaos, negative Lyapunov exponents approach zero. Immediately following the transition, at least six Lyapunov exponents are positive, indicating a high-dimensional chaotic attractor. Thus, high-dimensional chaos seems to set in directly, without going through a region of low-dimensional chaos.
@@
Figure S4: Multiple vortex solutions are possible at the same activity. a-d, Four representative snapshots of the stream function (color) and velocity field (arrows) overlaid on the nematic director (gray line integral convolution texture) at activity number $A=2400$. Each case results from a different seeding of random noise for the initial transient (see Methods).